OpenStudy (anonymous):

4t=48 ??? -x/3=27 ??? how do u solve these? alg 2

OpenStudy (anonymous):

Well fairly simple: \[t=\frac{ 48 }{ 4 } = 12\]

OpenStudy (anonymous):

thanks could u help with 2nd question in same message please???

OpenStudy (anonymous):

\[-x=27\times3\]\[-x=81\]\[x=-81\]

OpenStudy (anonymous):

14.)15-g=23-2g ???? I really don't know how to solve???

OpenStudy (anonymous):

\[15-g-15=23-2g-15\]\[-g = 8-2g\]\[-1g = 8 - 2g\] Same thing but I added a one so that you may understand. \[-1g + 2g = 8 - 2g +2g\]\[1g=8\] or \[g=8\]

OpenStudy (anonymous):

thanks soooo much(:

OpenStudy (anonymous):

Hopefully you understood.

OpenStudy (anonymous):

I did